OpenStudy (anonymous):

The position of an object at time t is given by s(t) = -2 - 6t. Find the instantaneous velocity at t = 2 by finding the derivative.

OpenStudy (shamim):

Tell me ds/dt=?

OpenStudy (anonymous):

Yah

OpenStudy (shamim):

U know d2/dt=0

OpenStudy (anonymous):

ok

OpenStudy (shamim):

Because derivative of any constant is =0

OpenStudy (anonymous):

ok

OpenStudy (shamim):

U know 2 is a number. So 2 is a constant. Right?

OpenStudy (anonymous):

yah 2 is a constant and thus the derivative is zero.. so would zero be the answer?

OpenStudy (shamim):

No

OpenStudy (anonymous):

This problem is just confusing me..

OpenStudy (shamim):

c d(6t)/dt=6*dt/dt=6*1=6

OpenStudy (anonymous):

I don't understand

OpenStudy (shamim):

ds/dt=d(2)/dt-d(6t)/dt

OpenStudy (shamim):

Nd velocity v=ds/dt

OpenStudy (anonymous):

ok... FLVS doesn't explain derivatives very well sorry

OpenStudy (shamim):

Plz tell me ur problem specifically

OpenStudy (shamim):

I wanna help u !

OpenStudy (shamim):

Dont u know Velocity v=ds/dt

OpenStudy (anonymous):

Thanks for the help.. What I think they want is a derivative question that involves solving with the t=2. This is why im confused

OpenStudy (anonymous):

I know that v=ds/dt

OpenStudy (shamim):

U know v=-6

OpenStudy (shamim):

So v is constant

OpenStudy (shamim):

After time t=2 sec the velocity is v=6

OpenStudy (shamim):

Nd it is negative

OpenStudy (anonymous):

ok... I think I was making it a lot more complicated than it rreally was

OpenStudy (shamim):

Negative sign indicates direction of velocity

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

Thanks!

OpenStudy (shamim):

U r most welcome !!

OpenStudy (shamim):

Anyway may i know the name of ur country?

OpenStudy (shamim):

Just for curiosity